OpenStudy (anonymous):

Evaluate the following integral. Integral of 6dx. -2 to 5

OpenStudy (anonymous):

\[\int\limits_{-2}^{5}6dx\]

OpenStudy (anonymous):

\[\int_{-2}^56~dx=6\left[x\right]_{-2}^5=6(5-(-2))\]